OpenStudy (anonymous):

what is the graph of 9x^2+16y^2=144?

OpenStudy (anonymous):

Do you only want answers ???

OpenStudy (anonymous):

@UnkleRhaukus wait..

OpenStudy (unklerhaukus):

\[9x^2+16y^2=144\] \[(3x)^2+(4y)^2=12^2\]

OpenStudy (anonymous):

This much is okay @UnkleRhaukus

OpenStudy (unklerhaukus):

does this remind you of a general form of an equation for some shape .?

OpenStudy (anonymous):

what is gthe axis of symmetry of this ellipse